# Logic Distillation: Learning from Code Function by Function for Planning and Decision-making

### v1

@@ Introduction
Figure 4: Function base. The L-LLM decomposes the rules and instantiates the planning and decision-making logic into multiple functions (each function performs a specific task, such as calculating distance, etc.). In addition, the L-LLM enables the S-LLM to accurately invoke relevant functions by creating a user manual (including rule explanations, function comments, corresponding invocation stages, etc.), thereby completing the planning and decision-making process.
@@
With Lagrange multiplier method, we get:
@@
[…]
Let Equation 9 be 00, we can get:
@@
Table 1: Results of pursuit game. In this context, “Success” refers to the successful conclusion of the game, where the blue dot captures the orange dot. “Failure without Violation” indicates an unsuccessful outcome due to the inability to capture the orange dot within the specified number of moves. On the other hand, “Failure with Violation” signifies an unsuccessful outcome resulting from a violation of the game rules. Lastly, “Average Steps of Success” quantifies the average number of moves required for successful completion of the game.
Success Failure without Violation Failure with Violation Average Steps of Success
S-LLM $0.00\%$ $0.00\%$ ${100.00\%}$ — steps
L-LLM $96.00\%$ $4.00\%$ ${0.00\%}$ 14.2214.22{14.22}14.22 steps
KD $88.50\%$ $10.50\%$ $1.00\%$ $15.23$ steps
[…]
(a) Failure of KD.
(b) Success of KD.
(c) Success of L-LLMs.
(d) LD.
Figure 5: The planning and decision-making outcome of one step in the pursuit game.
Experiments
Our experiments aim to: (1) verify the effectiveness of LD in planning and decision-making task, (2) compare LD with baselines from both global and local perspectives, and discuss the reasons for the effectiveness of LD (3) verify that LD has a stronger ability to respond to emergencies [^1] Code is included in the supplemental material and will be released upon the paper acceptance..
We conducted experiments based on the pursuit game to demonstrate that the proposed method can effectively enhance the planning and decision-making capabilities of S-LLMs. More specifically, the pursuit game involves two sides, each controlled by a different LLM. One LLM manages three blue dots, while the other LLM controls one orange dot. Each interaction between the two sides constitutes a step. In each iteration, the blue dots are constrained to move by two units, while the orange dot is restricted to a single unit of movement. The game concludes when the Manhattan distance between all three blue dots and the orange dot is less than 2 units.
In our experiments, the L-LLM is GLM-4, and the S-LLM is GLM4-9B-chat. The orange dot is consistently controlled by the L-LLM, while the blue dots are managed by the L-LLM, S-LLM, S-LLM with KD, and S-LLM with LD, respectively. Additional selection and judgment rules have been introduced to improve the success rate of the baselines. For selection, LLMs are provided with the next decision coordinates to choose from. Regarding judgment, if LLMs make more than seven illegal choices, the game is considered a failure. The upper limit for the number of moves in the game is capped at 100. We focus on utilizing LLMs to efficiently manage the blue dots to capture the orange dot. To perform KD, we initialize 221 sets of starting positions randomly and produce 103,355 sets of outputs with the L-LLM. Subsequently, we fine-tune the S-LLM with LoRA (Hu et al. 2021) based on these outputs. For a more details of the pursuit game, please refer to the Appendix.
For LD, as depicted in Figure 4, we initially establish a function base with the L-LLM, where L-LLM decomposes the rules and instantiating the planning and decision-making logic into multiple functions. Moreover, to assist the S-LLM in learning how to utilize various functions and decide when to invoke them, L-LLM creates a user manual for these functions. By fine-tuning the S-LLM with the user manual, it can comprehend the logic of the L-LLM and execute planning and decision-making processes function by function.
Better Performance with LD
The results of the pursuit game are presented in Table 1. It is evident that S-LLM struggles to comprehend complex instructions, as its failures consistently stem from rule violations. Conversely, the “Failure with Violation” rate of L-LLM is $0\%$, demonstrating its superior comprehension and caability to follow instructions. By employing KD to mimic the L-LLM’s outputs, the S-LLM’s planning and decision-making capabilities significantly improve. Nonetheless, its success rate is notably lower than that of the L-LLM, and the number of steps in successful instances is higher. As for LD, the game’s success rate has reached $100\%$, and the S-LLM completed the game with an average of $0.96$ fewer steps than the L-LLM. These results comprehensively demonstrate the effectiveness of LD in enhancing the S-LLM’s planning and decision-making capabilities.
@@
Figure 6: The overall process of the pursuit game. The trajectories of the three blue dots are represented by green, red, and grey dashed lines, respectively, with the darker colors indicating a higher number of passages.
Better Performance from a Local Perspective
We present the one-step planning and decision-making outcomes of different LLMs in Figure 5 (for the purpose of illustration, we have selected four scenarios where both the horizontal and vertical coordinates are within five units).
Figure 5(a) illustrates that while KD improves the alignment of S-LLM’s outputs with the rules, S-LLM still faces challenges in grasping the planning and decision-making logic of L-LLM. Consequently, the blue dots, controlled by S-LLM with KD, moves aimlessly, surpassing the maximum number of allowed movement steps, ultimately leading to the game’s failure. Figure 5(b) depicts that while S-LLM successfully completes the game, it still makes irrational decisions for the blue dots during specific moves, leading to a significantly higher number of steps compared to L-LLM. Regarding L-LLM and S-LLM with LD, as depicted in Figures 5(c) and 5(d), all LLMs demonstrate proficient control over the various blue dots, even as the blue dots approach the orange dots from different positions. These results indicate that, compared to KD, S-LLM with LD more effectively grasps the planning and decision-making logic of L-LLM.
@@
Success Failure without Violation Failure with Violation Average Steps of Success
S-LLM $0.00\%$ $0.00\%$ ${100.00\%}$ — steps
L-LLM $90.00\%$ $7.50\%$ ${2.50\%}$ $18.98$ steps
KD $52.00\%$ $2.50\%$ $45.50\%$ $18.5$ steps
[…]
Better Performance from a Global Perspective
@@
In Figure 6, we initialize different LLMs from identical starting positions: blue dots at $(3,8),(14,19),$, $(17,2)$, and the orange dot at $(20,18)$ to enable a global comparison of the overall planning and decision-making capabilities among different LLMs.
@@
In Figure 6(a), as S-LLM with KD merely mimics the outputs of L-LLM, blue dots may represent outputs from the L-LLM in different scenarios, resulting in behaviors such as repetitive circling. For instance, the point along the grey trajectory continuously shuttle back and forth, making it impossible to catch up with the orange point. In Figure 6(b), the point controlled by L-LLM appears to backtrack, mainly because of the orange point’s continuous back-and-forth movements in an attempt to escape encirclement. Besides, from Figure 6(c), it can be observed that S-LLM with LD enables the blue dots to approach the orange dot in a more direct manner. Contrasting with L-LLM, S-LLM with LD requires fewer steps to successfully capture the target. Such differences stem from different emphases: L-LLM employs a global perspective to design each function with rules. However, during each decision-making step, L-LLM was more susceptible to the influence of the current state, consequently neglecting global planning.
Pursuit Game with Emergencies
@@
In order to assess the capacity of different LLMs to handle emergencies, we introduced a $5\times 5$ restricted area within the game plane. Related results are presented in Table 2.
@@
In the more intricate scenario, L-LLM can still grasp the rules through simple textual descriptions and implement effective planning and decision-making, as the success rate achieve $90\%$. Conversely, S-LLM with KD achieves a success rate of $52\%$, with failures resulting from rule violations (accounting for $45.5\%$), indicating an inability of S-LLM with KD to comprehend the new rules. As for S-LLM with LD, the process involves S-LLM initially abstracting the restricted area as a coordinate filtering function. Then, this function will handle the output of $filter\_valid\_moves$ from Figure 4, producing a list that excludes the coordinates of the restricted area. Subsequently, this list is utilized by $select\_best\_move$ to generate a suitable coordinate. The success rate of S-LLM with LD exceeds that of L-LLM by $10\%$, and its average number of steps is approximately $4$ steps fewer than the baselines, demonstrating the powerful general capabilities of LD. It should be noted that, compared to KD, LD only uses few function usage examples to fine-tune S-LLM, which enable S-LLM to retain more general capabilities. More discussion please refer to the Appendix.
@@
In Figure 7, we illustrate the comprehensive planning and decision-making processes of L-LLM and S-LLM with LD, which further validates the effectiveness of converting logic into functions.
@@
Conclusion
LLMs have been widely applied across many different fields. However, larger LLMs (L-LLMs) with powerful capabilities in comprehension, planning, and decision-making are difficult to deploy on the vast majority of devices due to their parameter scale. In contrast to L-LLMs, smaller open-source LLMs (S-LLMs) are easier to deploy but fall significantly short in performance compared to their larger counterparts. To improve the performance of S-LLMs, researchers have proposed various Knowledge Distillation (KD) methods. Nevertheless, KD merely enables S-LLMs to mimic the outputs of L-LLMs, which is insufficient for addressing complex planning and decision-making problems. Thus, we propose Logic Distillation (LD), a method that instantiates the logic of L-LLMs by converting human-provided rules into functions, thereby establishing a function base. Subsequently, through fine-tuning, S-LLMs will comprehend the usage of each function, enabling S-LLMs to plan and make decisions function by function.

### v2

@@ 3.5 Why Selection Is Better
Figure 4: Function base. The L-LLM decomposes the rules and instantiates the decision-making logic into multiple functions (each function performs a specific task, such as calculating distance, etc.). In addition, the L-LLM enables the S-LLM to accurately invoke relevant functions by creating a user manual (including explanations, function comments, corresponding invocation stages, etc.), thereby completing the decision-making process.
@@
With Lagrange multiplier method Liu (1972), we get:
@@
$\left\{\begin{aligned} &\frac{\partial Q}{\partial p^{{}^{\prime}}_{\theta_{S}}(t_{i})}=-\log p^{{}^{\prime}}_{\theta_{S}}(t_{i})-1+\lambda\\ &\frac{\partial Q}{\partial\lambda}=\sum_{i=1}^{M}p^{{}^{\prime}}_{\theta_{S}}(t_{i})-1\end{aligned}\right$. (9)
Let Equation 9 be 0, we can get:
@@
Methods Success Failure without Violation Failure with Violation Average Steps of Success
GLM4 (Large) 96.00% 4.00% 0.00% 14.22 steps
GLM4-9B (Small) 0.00% 0.00% 100.00%
GLM4-9B-KD 88.50% 10.50% 1.00% 15.23 steps
GLM4-9B-LD 100.00% 0.00% 0.00% 13.26 steps
LLaMA3-70B (Large) 94.00% 2.00% 2.00% 16.37 steps
LLaMA3-7B (Small) 0.00% 2.00% 98.00%
LLaMA3-7B-KD 80.50% 11.50% 8.00% 18.74 steps
LLaMA3-7B-LD 100.00% 0.00% 0.00% 15.28 steps
Qwen2.5-72B (Large) 95.50% 4.00% 0.50% 13.92 steps
Qwen2.5-7B (Small) 1.00% 0.00% 99.00% 56.85 steps
Qwen2.5-7B-KD 89.00% 6.00% 5.00% 14.99 steps
Qwen2.5-7B-LD 100.00% 0.00% 0.00% 13.23 steps
Table 1: Results of pursuit game. In this context, “Success” refers to the successful conclusion of the game, where the blue dot captures the orange dot. “Failure without Violation” indicates an unsuccessful outcome due to the inability to capture the orange dot within the specified number of moves. On the other hand, “Failure with Violation” signifies an unsuccessful outcome resulting from a violation of the game rules. Lastly, “Average Steps of Success” quantifies the average number of moves required for successful completion of the game. All methods are tested on 200 sets of starting positions.
@@
Figure 5: The global perspective of the pursuit game based on GLM4 and GLM4-9B. The trajectories of the three blue dots are represented by green, red, and grey dashed lines, respectively, with the darker colors indicating a higher number of passages.
4 Experiments
Our experiments aim to: (1) verify the effectiveness of LD in decision-making tasks, (2) explore the reasons for the effectiveness of LD (3) verify that LD has a stronger ability to respond to emergencies. In our experiments, the L-LLMs are GLM-4, LLaMA3-70B and Qwen2.5-72B , while the S-LLMs are GLM4-9B, LLaMA3-7B and Qwen2.5-7B.
4.1 Better Performance in Pursuit Game
We first conduct experiments based on the pursuit game to demonstrate that the proposed method can effectively enhance the decision-making capabilities of S-LLMs. More specifically, the pursuit game involves two sides, each controlled by a different LLM. One LLM manages three blue dots, while the other one controls an orange dot. Each interaction between the two sides constitutes a step. In each iteration, the blue dots are constrained to move by two units, while the orange dot is restricted to a single unit of movement. The game concludes when the Manhattan distance between all three blue dots and the orange dot is less than 2 units.
Specifically, the orange dot is consistently controlled by the original S-LLM, while the blue dots are managed by the L-LLM, S-LLM, S-LLM with KD, and S-LLM with LD, respectively. Additional selection and judgment rules have been introduced to improve the success rate of the baselines. For selection, LLMs are provided with the next decision coordinates to choose from. Regarding judgment, if LLMs make more than seven illegal choices, the game is considered a failure. The upper limit for the number of moves in the game is capped at 100. To perform KD, we initialize 221 sets of starting positions randomly and produce 103,355 sets of outputs with the L-LLM. Subsequently, we fine-tune the S-LLM with LoRA Hu et al. (2021) based on these outputs. For LD, as depicted in Figure 4, we initially establish a function base with the L-LLM, where L-LLM decomposes the rules and instantiating the decision-making logic into multiple functions. Moreover, to assist the S-LLM in learning how to utilize various functions and decide when to invoke them, L-LLM creates a user manual for these functions. By fine-tuning S-LLM with the user manual, it can comprehend the logic of the L-LLM and execute decision-making processes function by function.
The results of the pursuit game are presented in Table 1. It is evident that S-LLM struggles to comprehend complex instructions, as its failures mainly stem from rule violations. Conversely, the “Failure with Violation” rates of L-LLMs are close to zero., demonstrating its superior comprehension and capability to follow instructions. By employing KD to mimic L-LLMs’ outputs, S-LLMs’ decision-making capabilities significantly improve. Nonetheless, their success rates are still notably lower than that of L-LLMs, and the number of steps in successful instances are higher. As for LD, the game’s success rates have reached $100\%$, and the average number of steps taken by S-LLMs is fewer than that of L-LLMs. These results comprehensively demonstrate the effectiveness of LD in enhancing S-LLMs’ decision-making capabilities.
Methods Success Failure without Violation Failure with Violation Average Steps of Success
GLM4 (Large) 90.00% 7.50% 2.50% 18.98 steps
GLM4-9B (Small) 0.00% 0.00% 100.00%
GLM4-9B-KD 52.00% 2.50% 45.50% 18.5 steps
GLM4-9B-LD 100.00% 0.00% 0.00% 14.65 steps
@@ 4.1 Better Performance in Pursuit Game
4.2 Why LD is Better
@@
In Figure 5, we initialize different LLMs from identical starting positions: blue dots at $(3,8),(14,19),(17,2)$, and the orange dot at $(20,18)$ to enable a global comparison of the overall decision-making capabilities among different LLMs.
@@
In Figure 5(a), as S-LLM with KD merely mimics the outputs of L-LLM, blue dots may represent outputs from the L-LLM in different scenarios, resulting in behaviors such as repetitive circling. For instance, the point along the grey trajectory continuously shuttle back and forth, making it impossible to catch up with the orange point. In Figure 5(b), the point controlled by L-LLM appears to backtrack, mainly because of the orange point’s continuous back-and-forth movements in an attempt to escape encirclement. Besides, from Figure 5(c), it can be observed that S-LLM with LD enables the blue dots to approach the orange dot in a more direct manner. Thus, contrasting with L-LLM, S-LLM with LD requires fewer steps to successfully capture the target. Such results stem from different emphases: For LD, L-LLM employs a global perspective to design functions, causing the S-LLM to pay more attention to the overall distance from the orange dot. However, during one decision-making step, L-LLM is more susceptible to the influence of the current state, consequently neglecting the global planning. This is also why the performance of LD in Table 1 is superior to that of L-LLM.
4.3 Pursuit Game with Emergencies
@@
In order to assess the capacity of different LLMs to handle emergencies, we introduced a $5\times 5$ restricted area within the game plane. Related results based on GLM4 and GLM4-9B are presented in Table 2.
@@
In the more intricate scenario, L-LLM can still grasp the rules through simple textual descriptions and implement effective decision-making, as the success rate achieve $90\%$. Conversely, S-LLM with KD achieves a success rate of $52\%$, with failures resulting from rule violations (accounting for $45.5\%$), indicating an inability of S-LLM with KD to comprehend the new rules. As for S-LLM with LD, the process involves S-LLM initially abstracting the restricted area as a coordinate filtering function. Then, this function will handle the output of filter_valid_movesfilter\_valid\_moves from Figure 4, producing a list that excludes the coordinates of the restricted area. Subsequently, this list is utilized by select_best_moveselect\_best\_move to generate a suitable coordinate. The success rate of S-LLM with LD exceeds that of L-LLM by $10\%$, and its average number of steps is approximately $4$ steps fewer than the baselines, demonstrating the powerful general capabilities of LD. It should be noted that, compared to KD, LD only uses a few functional examples to fine-tune the S-LLM, enabling S-LLM to retain more general capabilities.
@@
In Figure 6, we illustrate the comprehensive decision-making processes of L-LLM and S-LLM with LD, which further validates LD has a stronger ability to respond to emergencies.
This experiment also highlights the advantages of controlling embodied agents through LLMs. Traditional reinforcement learning methods for controlling embodied agents require retraining the model from scratch when encountering new influencing factors. In contrast, under the LD framework, S-LLMs can directly accept new rules and add them to the function library, enabling the agent to address new challenges by combining different functions.
@@
4.4 Better Performance in 21 Ponits
To further validate the improvement in S-LLM’s decision-making capabilities with LD, we engage LLMs in a modified version of the game “21 points.” As depicted in Figure 7, during each round, the two participating LLMs must decide whether to “stand” or “hit”, with the objective of surpassing their opponent’s total score without exceeding 21 points. Notably, cards “10”, “J”, “Q”, and “K” are each assigned a value of 10, and card “A” can be used as either 1 or 11. Furthermore, we streamline the game by reducing the deck to 26 cards of two suits, thereby enabling the LLMs to deduce the likelihood of exceeding 21 points after requesting a card, based on the available information.
Methods Victory Loss Draw
GLM4 (Large) 55.50% 42.00% 2.50%
GLM4-9B (Small) 36.50% 62.00% 1.50%
GLM4-9B-KD 48.00% 49.00% 3.00%
GLM4-9B-LD 59.50% 38.50% 2.00%
LLaMA3-70B (Large) 57.00% 42.00% 1.00%
LLaMA3-7B (Small) 39.50% 60.00% 0.50%
LLaMA3-7B-KD 48.50% 58.00% 3.50%
LLaMA3-7B-LD 58.50% 38.50% 3.00%
Qwen2.5-72B (Large) 50.50% 46.50% 3.00%
Qwen2.5-7B (Small) 34.00% 65.50% 0.50%
Qwen2.5-7B-KD 44.50% 54.50% 0.00%
Qwen2.5-7B-LD 55.00% 43.00% 2.00%
Table 3: The victories, losses, and draws of each LLM in 200 times of 21 points.
Related results are presented in Table 3. It can be observed that L-LLMs demonstrate significant advantages in decision-making, as their win rates substantially exceed their rates of losses and draws. As for S-LLMs with KD, we first fine-tune S-LLMs with the outputs of L-LLMs. Although KD enhances S-LLMs’ decision-making capabilities, the improvement is not significant, and there is still a gap between S-LLMs and L-LLMs. In contrast, S-LLMs with LD achieve comparable or even superior results compaed to L-LLMs. This is because L-LLMs teach the logic of decision-making to S-LLMs in the form of functions, enabling S-LLMs to maintain a global perspective. However, L-LLMs are susceptible to the influence of the current state in the decision-making process at each step, thereby losing its ability for global planning.
5 Conclusion
LLMs have been widely applied across many different fields. However, larger LLMs (L-LLMs) with powerful capabilities are difficult to deploy on the vast majority of devices due to their parameter scale. In contrast to L-LLMs, smaller open-source LLMs (S-LLMs) are easier to deploy but fall significantly short in performance compared to their larger counterparts. To improve the performance of S-LLMs, researchers have proposed various Knowledge Distillation (KD) methods. Nevertheless, KD merely enables S-LLMs to mimic the outputs of L-LLMs, which is insufficient for addressing decision-making problems. Thus, we propose Logic Distillation (LD), a method that instantiates the logic of L-LLMs by converting instructions into functions, thereby establishing a function base. Subsequently, through fine-tuning, S-LLMs will comprehend the usage of each function, enabling S-LLMs to make decisions. Experimental results demonstrate that the proposed method, with a small amount of example fine-tuning, can enable S-LLMs to match or even surpass the decision-making capabilities of L-LLMs.
Acknowledgments
This work is supported by National Natural Science Foundation of China (Grant Nos. 62436007, 62325602, 62036010, 62276238, U24A20326), the Foundation of Henan Educational Committee under Grant (No.25HASTIT034), the Natural Science Foundation of Henan under Grant (No.232300421095), the Key R&D Projects in Zhejiang Province (No.2024C01106, 2025C01030), the Zhejiang NSF (LRG25F020001).